OpenStudy (anonymous):

How many grams of magnesium phosphate are needed to make 250mls of a .25 molar solution

OpenStudy (anonymous):

16,4 g

OpenStudy (xishem):

Molar concentration is modeled by the following equation...\[c_i=\frac{n_i}{V_{sol'n}}\]Where... c_i is the molar concentration of the solution in substance i n_i is the number of mols of i in the solution V_sol'n is the total volume of the solution Of those three variables, we know the total volume of the solution and molar concentration. Let's go ahead and express the number of moles in terms of mass for magnesium phosphate... \[MM_i=\frac{m_i}{n_i} \rightarrow n_i=\frac{m_i}{MM_i}\]\[MM_{Mg_3(PO_4)_2}=262.85\frac{g}{mol}\] And now...\[c_i=\frac{n_i}{V_{sol'n}} \rightarrow c_i=\frac{\frac{m_i}{MM_i}}{V_{sol'n}} \rightarrow m_i=(c_i)(V_{sol'n})(MM_i)\]\[m_i=(0.25\frac{mol}{L})(0.250L)(262.85\frac{g}{mol})=16.428\]\[m=16g\]